# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 3.6. Equidistribution within small interval
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (3.53)
We first obtain some estimate on the first integral of (3.6). The idea behind estimating this integral is recursive uses of change of variables that lead to consistent cancellation. We first do the change of variable $z_{m-2}\mapsto z_{m-2}+(y-\sum_{i=1}^{d-3}z_{m+i-d})/5$. We have
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}e^{-\frac{\sum_{i=1}^{d-3}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m-2}+\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}$
$\quad\quad\quad\cdot e^{-\frac{\frac{1}{4}\left(-z_{m-2}+4\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}dz_{m-2}\cdots dz_{m+1-d}.$ (3.54)
Notice that in the expansion of the exponents, we get a cancellation of the term $2z_{m-2}((y-\sum_{i=1}^{d-3}z_{m+i-d})/5)$. Hence, (3.6) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{\frac{5}{4}z_{m-2}^{2}}{2}}dz_{m-2}\right)dz_{m-3}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}dz_{m-3}\cdots dz_{m+1-d}$ (3.55)
We see that the estimate in (3.6) has a similar structure, which allows us to use induction to estimate the interval. For an arbitrary $k$ such that $1\leq k\leq d-2$, we want to look at
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k}(y-\sum_{i=1}^{k}z_{m+i-d})^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.56)
where $A_{k}>0$. Using the change of variable $z_{m+k-d}\mapsto z_{m+k-d}+\frac{A_{k}}{1+A_{k}}(y-\sum_{i=1}^{k-1}z_{m+i-d})$, (3.56) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m+k-d}+\frac{A_{k}}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}$
$\quad\quad\quad\cdot e^{-\frac{A_{k}\left(-z_{m+k-d}+\frac{1}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d}.$ (3.57)
Expanding the exponents also gives a cancellation of the term
$\frac{2A_{k}}{1+A_{k}}z_{m+k-d}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right).$ (3.58)
Hence, (3.56) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}e^{-\frac{(1+A_{k})z^{2}_{m+k-d}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.59)
where
$A_{k-1}\ :=\ \frac{A_{k}^{2}}{1+A_{k}^{2}}+\frac{A_{k}}{(1+A_{k})^{2}}.$ (3.60)
Note that since $A_{k}>0$, then $A_{k+1}>0$. Hence, (3.56) becomes bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{(1+A_{k})z_{m+k-d}^{2}}{2}}dz_{m+k-d}\right)$
$\quad\quad\quad dz_{m+(k-1)-d}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}dz_{m+(k-1)-d}\cdots dz_{m+1-d},$ (3.61)
Thus, by induction, (3.6) becomes bounded above by
$e^{-(A_{0}y^{2})/2},$ (3.62)
where $A_{0}>0$ is recursively defined by (3.60) with $A_{d-2}=1/4$. Similarly, we can show by induction that the second integral in (3.6) has the estimate
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right)dz_{m-2}\cdots dz_{m+1-d}\ \ll\ e^{-(B_{0}y^{2})/2}$ (3.63)
for some $B_{0}>0$. Hence, if $d>2$, (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ e^{-(A_{0}y^{2})/2}+e^{-(B_{0}y^{2})/2}\ \ll\ e^{-(C_{0}y^{2})/2},$ (3.64)
where $C_{0}:=\min\{A_{0},B_{0}\}$. Recall from (3.6) that when $d=2$,
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right).$ (3.65)
@@
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}y^{2}}{2}}+e^{-\frac{D_{0}\left(y+C\sqrt{N}\right)^{2}}{2}},$ (3.66)
where $D_{0}:=\min\{C_{0},1/4\}$. Note that $e^{-D_{0}y^{2}/2}$ and $\varphi(\sqrt{D_{0}}(y+C\sqrt{N}))$ each has only one global extreme, at $y=0$ and $y=-C\sqrt{N}$ respectively. Hence, locally on $n/\sqrt{N},(n+1)/\sqrt{N}]$ the functions are monotonic and can be bounded above by the sum of its values at the two end points. Hence, for all $n$,
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}},$ (3.67)
for all $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]$. Thus, returning to (3.45), for all $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$,
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|$
$\quad\ \ll\ \frac{1}{\sqrt{N}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.68)
@@
$e_{n}(a,b)\ \leq\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|dy$
$\quad\ \ll\ \left(\frac{1}{\sqrt{N}}\right)^{2}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.69)
3.7. Upper bound on error term $\mathcal{M}_{\textup{err},N}(a,b)$
@@
$\mathcal{M}_{\textup{err,N}}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.70)
Based on the estimate $e_{n}(a,b)$ in (3.6), we can pull out one of the $1/\sqrt{N}$ factors in the bracket and obtain
$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)$
$\ \ll\ \frac{1}{\sqrt{N}}\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right)\right).$ (3.71)
We observe that the above in the parentheses are Riemann sums for $e^{-D_{0}x^{2}/2}$ and $e^{-D_{0}(x+C\sqrt{N})^{2}/2}$ on $(-\infty,\infty)$, which are both finite. Hence
$\mathcal{M}_{\textup{err},N}(a,b)\ \ll\ \frac{1}{\sqrt{N}}.$ (3.72)
@@
3.8. Evaluation of main term $\mathcal{M}_{\textup{main},N}(a,b)$
Finally, we establish the main term $\mathcal{M}_{\textup{main},N}(a,b)$ of the main term $\mathcal{M}_{N}(a,b)$. From (3.37) and (3.42), we see that
$\mathcal{M}_{\textup{main},N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$
$\ =\ (b-a)\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right).$ (3.73)
Since the term in the parentheses above is a Riemann sum for $\int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy$, we have
$\frac{1}{\sqrt{N}}\sum_{n=-2C\cdot N}^{2C\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)$ $\ =\ \int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy+o(1)$
$\ =\ \mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty)+o(1).$ (3.74)
Hence the main term $\mathcal{M}_{\textup{main},N}(a,b)$ becomes
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)(\mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty))+o(1).$ (3.75)
@@ 3.8. Evaluation of main term $\mathcal{M}_{\textup{main},N}(a,b)$
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.76)
To simplify our calculation, we want to extend the interval of integration of the outer integral to -$\infty$. We do so by showing that the tail of the integral is small, i.e.,
$T_{N}\ :=\ \int_{-\infty}^{-C\sqrt{N}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}\ \approx\ 0.$ (3.77)
We have
$T_{N}$ $\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}(\Phi(z_{m+1-d}))^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})dz_{m}\cdots dz_{m+1-d}$
$\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\varphi(z_{m+1-d})dz_{m+1-d},$ (3.78)
Hence, we have reduced the problem to showing the smallness of the Gaussian tail. We first make a definition.
Definition 3.5.
For functions $f(x)$ and $g(x)$, we say that $f(x)=\Theta(g(x))$ if $f(x)=O(g(x))$ and $g(x)=O(f(x))$.
The following result is a straightforward calculation that provides an upper bound on the Gaussian tail.
Proposition 3.6.
[Mil1] Suppose that $g:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is a function such that $g(N)=\Theta(N^{\epsilon^{\prime}})$ for some $\epsilon^{\prime}>0$. Then
$\int_{-\infty}^{-g(N)}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ =\ \int_{g(N)}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ \ll\ e^{-g(N)/2}.$ (3.79)
Back to (3.8), by Proposition 3.6
$T_{N}\ \ll\ e^{-C\sqrt{N}}.$ (3.80)
Thus, (3.76) becomes
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}+O\left(e^{-C\sqrt{N}}\right).$ (3.81)
Now, by Proposition 3.3,
$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ (3.82)
is the joint PDF of $W_{(m+1-d)},\dots,W_{(m)}$, where $W_{1},\dots,W_{m}$ are i.i.d. $\sim N(0,1)$. Thus, following our derivation in Subsection 3.2,
$H(y)\ :=\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.83)
is the CDF of $\sum_{i=1}^{d}W_{(m+i-d)}$. We know that for a cumulative density function $H(y)$, $H(-\infty)=0$ and $H(\infty)=1$. Hence,
$\mathcal{M}_{Y^{(N)}}(-\infty)$ $\ =\ H(-\infty)+O(e^{-C\sqrt{N}})\ =\ O\left(e^{-C\sqrt{N}}\right)$
$\mathcal{M}_{Y^{(N)}}(\infty)$ $\ =\ H(\infty)+O(e^{-C\sqrt{N}})\ =\ 1+O\left(e^{-C\sqrt{N}}\right).$ (3.84)
Thus, substituting these two estimates into (3.75) yields
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)\left(1+O\left(e^{-C\sqrt{N}}\right)\right)+o(1))\ =\ (b-a)+o(1).$ (3.85)
Returning to (3.39) and (3.38), combined with the estimate for the error term $\mathcal{E}_{N}(a,b)$ in (3.4) and the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of $\mathcal{M}_{N}(a,b)$ in (3.72), we have
@@
$\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)+\mathcal{E}_{N}(a,b)$
$\ =\ (b-a)+o(1)+O\left(\frac{1}{\sqrt{N}}\right)+O(N^{-\delta})$
$\ =\ (b-a)+o(1).$ (3.86)
@@
Appendix A Proof of Proposition 2.5
In this section, we prove Proposition 2.5 which provides a quantitative bound on the difference among probabilities within an interval. Let us first recall the statement of Proposition 2.5.
Proposition 2.5.

### v3

@@ 3.6 Equidistribution within small interval
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (134)
Notice that the first integral in (3.6) is the PDF of $W_{1}+\cdots+W_{d-2}+2W_{d-1}$, where $W_{i}$’s are i.i.d. standard Gaussians, and the second integral in (3.6) is the PDF of $V_{1}+\cdots+V_{d-2}+3/2V_{d-1}$, where the $V_{i}$’s are also i.i.d. standard Gaussians. Hence, there exists a constant $C_{0}>0$ such that
$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-(C_{0}y^{2})/2}.$ (135)
Moreover, recall from (3.6) that when $d=2$,
$\frac{d}{dy}I_{N}(y)\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+N^{\delta^{\prime}}\right).$ (136)
@@
$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-\frac{D_{0}y^{2}}{2}}+e^{-\frac{D_{0}\left(y+N^{\delta^{\prime}}\right)^{2}}{2}},$ (137)
where $D_{0}:=\min\{C_{0},1/4\}$. Note that $e^{-D_{0}y^{2}/2}$ and $\varphi(\sqrt{D_{0}}(y+N^{\delta^{\prime}}))$ each has only one global extreme, at $y=0$ and $y=-N^{\delta^{\prime}}$ respectively. Hence, locally on $n/\sqrt{N},(n+1)/\sqrt{N}]$ the functions are monotonic and can be bounded above by the sum of its values at the two end points. Hence, for all $n$,
$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}},$ (138)
for all $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]$. Now, we return to (3.6). For all $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$,
$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{1})-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{2})\right|$
$\quad\ \ll\ \frac{1}{\sqrt{N}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)+O(N^{-2\delta^{\prime}}).$ (139)
@@
$e_{n}(a,b)$ $\ \leq\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right|dy$
$\ \ll\ \left(\frac{1}{\sqrt{N}}\right)^{2}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)$
$\quad\quad+O(N^{-1/2-2\delta^{\prime}}).$ (140)
3.7 Upper bound on error term $\mathcal{M}_{N,\textup{ error}}(a,b)$
@@
$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)\ \approx\ 0.$ (141)
Based on the estimate of $e_{n}(a,b)$ in (3.6), we can pull out one of the $1/\sqrt{N}$ factors and obtain
$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)$
$\ \ll\ \frac{1}{\sqrt{N}}\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)\right)$
$\quad\quad+O(N^{-\delta^{\prime}}).$ (142)
We observe that item in the parentheses above is double of the sum of the Riemann sums for $\int_{-\infty}^{\infty}e^{-D_{0}x^{2}/2}dx$ and $\int_{-\infty}^{\infty}e^{-D_{0}(x+N^{\delta^{\prime}})^{2}/2}dx$, which are both finite. Hence,
[…]
@@
3.8 Evaluation of main term $\mathcal{M}_{N,\textup{ main}}(a,b)$
Finally, we establish the main term $\mathcal{M}_{N,\textup{ main}}(a,b)$ of $\mathcal{M}_{N}(a,b)$. From (117) and (122), we see that
$\mathcal{M}_{N,\textup{ main}}(a,b)$ $\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}m_{n}(a,b)$
$\ =\ (b-a)\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right).$ (144)
Since the item in the parentheses above is the Riemann sum for $\int_{-\infty}^{\infty}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy$, we have
$\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)$ $\ =\ \int_{-\infty}^{\infty}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy+o(1)$
$\ =\ F_{Y^{(N)},\textup{ main}}(\infty)-F_{Y^{(N)},\textup{ main}}(-\infty)+o(1).$ (145)
Hence the main term $\mathcal{M}_{N,\textup{ main}}(a,b)$ becomes
$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ (b-a)(F_{Y^{(N)},\textup{ main}}(\infty)-F_{Y^{(N),\textup{ main}}}(-\infty))+o(1).$ (146)
@@ 3.8 Evaluation of main term $\mathcal{M}_{N,\textup{ main}}(a,b)$
$F_{Y^{(N)},\textup{ main}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (147)
Now, by Proposition 5 and (81),
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ (148)
is the joint PDF of $W_{(m+1-d)},\dots,W_{(m)}$, where $W_{1},\dots,W_{m}$ are i.i.d. standard Gaussians. Thus, following our derivation in Subsection 3.2,
$F_{Y^{(N)},\textup{ main}}(y)\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (149)
is the CDF of $\sum_{i=1}^{d}W_{(m+i-d)}$. We know that for a cumulative density function $F(y)$, $F(-\infty)=0$ and $F(\infty)=1$. Thus, substituting these two estimates into (146) yields
$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ (b-a)+o(1).$ (150)
Returning to (119) and (118), combined with the estimate for the error term $\mathcal{E}_{N}(a,b)$ in (3.4) and the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of $\mathcal{M}_{N}(a,b)$ in (143), we have
@@
[…]
$\ =\ (b-a)+o(1)+O(N^{-\delta^{\prime}})+O(N^{-\delta^{\prime}})$
$\ =\ (b-a)+o(1).$ (151)
@@
4 Funding
This work was partially supported by Williams College Summer Science Program Research Fellowship, the Finnerty Fund, and NSF Grant DMS2241623. We would like to thank the anonymous referee for their valuable time and constructive comments.
Appendix A Proof of Proposition 3
In this section, we prove Proposition 3, which provides a quantitative bound on the difference among probabilities within an interval. Let us first recall the statement of Proposition 3.
Proposition 3.